# Inertial Coordination Games

### v1

@@ Abstract
We analyze inertial coordination games: dynamic coordination games with an endogenously changing state that depends on (i) a persistent fundamental that players privately learn about; and (ii) past play. We give a tight characterization of how the speed of learning shapes equilibrium dynamics: the risk-dominant action is selected in the limit if and only if learning is slow such that posterior precisions grow sub-quadratically. This generalizes results from static global games and endows them with an alternate learning foundation. Conversely, when learning is fast, equilibrium dynamics exhibit persistence and limit play is shaped by initial play. Whenever the risk dominant equilibrium is selected, the path of play undergoes a sudden transition when signals are precise, and a gradual transition when signals are noisy.
@@ 1   Introduction
Coordination games are at the heart of many economic phenomena. In such games, shocks to the fundamentals can sometimes propagate: as traders attack a currency peg, the central bank has to expand reserves to maintain the peg; as reserves are drained, more traders are incentivized to take advantage of this weakness which, in turn, further drains the central bank’s reserves, enticing even more attacks. But shocks can also fizzle out: traders might learn quite quickly that the central bank’s balance sheet is healthy so that the proportion of the population who are (incorrectly) pessimistic thins out rapidly. Thus, this downward spiral of an initial attack begetting further attacks eventually loses steam. [^1] Similar forces operate with bank runs (via bank liquidity), networked products (via product improvements), firm pricing (via sticky consumers). When do shocks propagate and become self-fulfilling, and when do they fizzle out?
We will show that the speed of private learning is a crucial determinant of whether shocks exhibit persistence. Our model is a dynamic coordination game with an endogenously changing state. A continuum of agents repeatedly decide whether or not to take a risky action whose payoff comprises (i) a persistent fundamental component; and (ii) an endogeneous component which depends on past play. In each period, each agent observes a private signal about the fundamental and makes inferences about the current state, which determines her current behavior. Agents’ current aggregate decisions, in turn, propagate into the future by shaping the evolution of future states. Thus, our environment features both private learning—in which dispersion in beliefs about fundamentals across the population shrinks with time—as well as intertemporal contagion, in which higher current play increases future states which incentivizes higher future play. This endogenous inertia in how the state evolves plays a critical role in the dynamics of play and we call such environments inertial coordination games. [^2] Our model is consistent with extensive empirical documentation of bank failures: Correia, Luck, and Verner (2024) find that “most bank failures are the result of a deterioration in bank solvency.” (in our model, a shock) rather than being purely self-fulfilling. Furthermore, they find that “ Depositors tend to be slow to react to an increased risk of bank failure” (in our model, the intertia).
@@
Our main result (Theorem 1) develops a tight characterization of the relationship between the speed of learning and limit play. We identify a critical threshold for the speed of learning that determines limit play: the risk-dominant action is played in the limit if and only if posterior precisions grow sub-quadratically. Sub-quadratic learning includes canonical learning environments such as (i) independent and identically distributed signals; and (ii) the receipt of a one-time signal. Super-quadratic learning includes social learning, where in each period, agents are randomly matched and sample their partner’s information sets.
@@
Indeed, static global games (Carlsson and Van Damme, 1993) are recovered exactly as a special case of our model in which players only receive a single signal at time $t=1$: since this learning process is sub-quadratic (posterior precisions are unchanged after $t=1$), the risk-dominant profile is—as in global games—played in the limit. This connection arises from observing that best-responding to expected past play along each time step is equivalent in our model to the interim elimination of strictly dominated strategies in the static game. As we make precise in Appendix D, this uncovers a formal connection between “contagion” arguments over type spaces (which encode hierarchies of higher-order beliefs) and “contagion” over time (where the current state depends on past play). Thus, our results generalize the predictions of static global games to sub-quadratic learning and offer a simple dynamic learning foundation for the ‘vanishing noise limit’ commonly invoked in that literature.
A notable implication of our main result is that slow learning implies history independence: the initial shock to aggregate play as well as any finite path of play are irrelevant for limit play. Conversely, when the posterior precision grows super-quadratically, learning is sufficiently fast which induces hysteresis and history-dependence that leads to non-risk dominant limit play. We further characterize all possible equilibria: any monotone selection between initial play-dominance (under common-knowledge of the fundamental state, players best-respond to the conjecture that the proportion of players choosing the risky action is given by initial play) and risk-dominance can be implemented via some appropriate learning rate. Thus, our results offer a simple and unified analysis for when our model’s dynamics exhibit history-dependence such that initial shocks can be self-fulfilling, [^3] See Section I of Krugman (1991) for an excellent discussion of history-dependence and expectations in economics. and when it exhibits history-independence so that limit play is independent of the initial shock.
@@
Finally, our model makes sharp qualitative predictions about the path of play (Proposition 3) whenever learning is slow enough such that the risk-dominant profile is played in the limit. In particular, transition dynamics toward limit play depend starkly on the precision of signals. When signals are precise, aggregate play exhibits a sudden transition from approximately all players playing the non-risk-dominant action to approximately all players playing the risk-dominant action. Conversely, when signals are noisy, aggregate play exhibits a gradual transition as the measure of players playing risk-dominant action increases gradually. These differing transition regimes are driven by differences in time-varying heterogeneity in beliefs. The importance of heterogeneity has been recognized as a crucial determinant of equilibrium uniqueness in coordination games (Morris and Shin, 2006). Our results complement this insight by highlighting that the rate of decrease of heterogeneity driven by learning is also a crucial determinant of the path of play. It also suggests that “spikes” in the time-series of aggregate actions (e.g., withdrawal decisions or adoption of networked goods) can be consistent with transition to limit equilibrium play and need not necessarily correspond to “equilibrium shifts” (Morris and Yildiz (2019)).
@@
Related literature. Our paper directly relates to the large literature on global games. A broad takeaway from the literature on static global games (Carlsson and Van Damme, 1993; Kajii and Morris, 1997; Frankel, Morris, and Pauzner, 2003, among others) is that perturbing common knowledge by introducing a vanishing amount of private uncertainty often selects the risk dominant equilibrium. [^4] But not always (Weinstein and Yildiz, 2007). This approach has been widely deployed in a variety of economic applications to select among multiple equilibria (see, e.g., Morris and Shin (1998); Goldstein and Pauzner (2005) among many others). Our model nests the canonical global games model of Carlsson and Van Damme (1993) with Normal signals as a special case with a private signal at $t=1$ and no learning thereafter. The “contagion argument” in global games where players iteratively eliminate interim dominated strategies is sometimes viewed metaphorically or as a convenient way to select among equilibria. [^5] See, Morris (2002) and Chapter 2 of Spiegler (2024). By developing a simple dynamic environment with endogeneous time-varying states, we thus offer an explicit dynamic foundation for the “contagion” logic of global games. Moreover, we generalize the prediction of risk-dominant selection to the regime of sub-quadratic learning.
@@
Our work also relates to dynamic global games (Angeletos, Hellwig, and Pavan, 2007; Dasgupta, 2007; Dasgupta, Steiner, and Stewart, 2012, among others). In much of this literature, coordination is contemporaneous i.e., time-$t$ play matters for time-$t$ incentives. [^7] Important exceptions are the papers of Dasgupta, Steiner, and Stewart (2012) and Mathevet and Steiner (2013). Our setting is one in which incentives are intertemporally linked because the path of past play shapes the current state which in turn shapes current incentives. In this regard, our paper is perhaps most related to the elegant paper of Mathevet and Steiner (2013) who study finite-time horizon dynamic global games with non-contemporaneous coordination motives where the terminal outcome depends on both the state and path of play. Our models are non-nested and deliver complementary insights. In our model, the state is time-varying and endogenous to past play so players’ payoffs depend on interim play which in turn drives the state-evolution. In Mathevet and Steiner (2013)’s model, ex-post payoffs depend on a binary outcome which is determined by a persistent fundamental and the path of aggregate play. This mapping from paths to a binary outcome allows Mathevet and Steiner (2013) to tractably handle a wide class of payoffs. By contrast, our payoffs are more specific, but evolve endogeneously and take on a continuum of values. We exploit the structure of our environment to derive sharp implications of learning speeds on limit outcomes. [^8] The main result of Mathevet and Steiner (2013) is that aggregate play at the critical state at which outcomes hinge only depends on payoffs; by contrast, we show that limit aggregate play at any state depends sharply on the speed of learning. Note also that Mathevet and Steiner (2013)’s definition of “fast learning” takes the precision of each signal arbitrarily large, which is sufficient for equilibrium existence. This is different from our notion which requires the (weaker) condition of super-quadratic growth in posterior precisions, and where limit equilibrium play always converges.
@@
Finally, our results are closely connected to the literature on learning in coordination games. One interpretation of our model is one of best-response dynamics with Bayesian learning about a fundamental state or past play. To our knowledge, this is novel to the extant literature on learning in games which assumes that past play is perfectly observed and there is no fundamental uncertainty. [^9] See Fudenberg and Levine (1998) for a survey. Our paper is related to an important paper of Crawford (1995) who studies adaptive dynamics in complete information coordination games by modelling player strategies as a linear adjustment rule to be estimated; this is similar to the interpretation of players in our model as best-responding to past play. Thus, a contribution of our paper is to analyze how Bayesian learning about fundamental uncertainty can interact with adaptive learning about strategic uncertainty (Camerer and Hua Ho, 1999). [^10] Camerer and Hua Ho (1999) develop a model of learning in complete information games which features both modes of learning; by contrast, in the best-response interpretation of our model, different modes of learning correspond to different kinds of uncertainty:‘fundamental uncertainty’ is resolved via Bayesian learning but ‘strategic uncertainty’ is resolved via adaptive or reinforcement learning.
@@ 2   Model
There is an unknown and persistent fundamental $\theta\in\mathbb{R}$. For each $t\in\mathcal{T}$, let $\theta_{t}$ be the state at time $t$ and for each $t>1$, [^12] For any $\alpha\in(0,1)$, analogs of our results apply for $\theta_{t}=\alpha\theta+(1-\alpha)\lambda_{t-1}$; see Appendix E.
@@
$\lambda_{0}$ is common knowledge but interim payoffs as well as $(\lambda_{t})_{t\geq 1}$ are unobserved and inferred via beliefs about the state $\theta$ [^13] An identical assumption is made by Angeletos, Hellwig, and Pavan (2007). Our model is equivalent to one in which agents observe noisy signals of past play $(\lambda_{t})_{t\geq 1}$, or both signals about the state as well as past play—we formalize this an equivalence in Appendix B and discuss this below.
Interpreting initial play. We interpret $\lambda_{0}\in[0,1]$ as a shock. This is equivalent to a one-off and transitory shock to the fundamental or costs at time $1$. [^14] More explicitly, suppose there is zero initial play but there is an initial shock to costs of playing the risky action. Costs are time-varying $c_{1},c_{2},\ldots$ with $c:=c_{2}=c_{3}=c_{4}=\ldots$ and $c_{1}\neq c$. Then, the dynamics of this model with zero initial play is equivalent to the model with $\lambda_{0}=c-c_{1}$. The impact of this shock is equivalent to that of non-zero initial play: positive shocks which incentive the risky action corresponds to high $\lambda_{0}$ and vice versa. In the language of macroeconomics, this corresponds to an unexpected shock (“MIT shock”) and we analyze the dynamics of aggregate play following the shock. In the language of fictitious play, this corresponds to a shock to initial play and we analyze how the speed of learning shapes the limit action profile.
@@
Limits and Learning. Let the learning process $\Sigma:=(\sigma_{t}^{2})_{t\in\mathcal{T}}$ be a strictly positive sequence of signal variances determining the speed of learning. Note that instead of $\Sigma$, we may equivalently directly specify a nonnegative, increasing sequence of posterior precisions $(\eta_{t}^{-2})_{t}$, since such a sequence is induced by $\sigma_{t}^{-2}:=\eta_{t}^{-2}-\eta_{t-1}^{-2}$. When convenient, we will therefore wlog consider $(\eta_{t}^{-2})_{t}$ as a primitive.
@@
It will be helpful to discuss various aspects of our model:
  • Discussion of state transition. Our model is one in which the current state is shaped by past play. For instance, in a currency crisis, if more players attacked the currency in the previous period, this weakens the fundamentals—e.g., the quantity of central bank reserves—which leads to a higher state today. With networked goods, if more players bought the good the previous period, this generates revenue or data for the firm to improve its product which leads to a better product today. We further note that although the current state $\theta_{t}$ only depends directly on $\lambda_{t-1}$, it also depends indirectly on the full (endogeneous) path of $(\lambda_{s})_{s<t-1}$; this is broadly analogous to how an AR$(1)$ process depends indirectly on realizations of the process which are far into the past (although our setting is non-stationary). [^18] An alternate interpretation is that our model is a form of best-response dynamics. Of course, some care is required with this interpretation since past play is unobserved: players are thus best-responding to their best guess of past play. Experimental evidence in (complete information) coordination games find that the vast majority (96%) of play constitute myopic best-responses (Mäs and Nax, 2016).
  • Discussion of Gaussian learning structure. We have explicitly parameterized the model with Gaussian signals; doing so allows us to analyze the law of motion of aggregate play as a discrete dynamical system. It will turn out, however, that our main result (Theorem 1) does not rely on the parametric form of Gaussian learning because we will show that whether risk dominance obtains is a tail event (Corollary 1) and, furthermore, under regularity assumptions, the posterior distribution generated from a sequence of bounded-variance noise structures converges to the normal distribution. [^19] By a straightforward application of the Bernstein–von Mises theorem (Van der Vaart, 2000). Thus, our result characterizing sub/super-quadratic boundary as critical to limit play holds for more general learning processes.
@@
  • Discussion of observability. We have assumed agents do not observe aggregate past play. [^20] Aggregate past play is an injective function of the state, so if agents were to observe this, they would immediately learn the state. However, our model is equivalent to one in which agents receive private signals both about the state, and about aggregate past play. [^21] For instance, see Dasgupta (2007); Trevino (2020). Explicitly, suppose that agent $i\in[0,1]$, in addition to observing $x_{it}\sim N(\theta,\sigma^{2}_{t})$, additionally observes an independent signal distributed $y_{it}\sim N(\Phi^{-1}(\lambda_{t-1}),\tau_{t}^{2})$; this signal structure was introduced by Dasgupta (2007) and used recently by Trevino (2020) to study financial contagion. Then, we can show by induction that equilibrium behavior is equivalent to a model in which players receive a more precise signal about $\theta$ and no information about aggregate past play; Appendix B formalizes this connection. In light of this observation, we load all learning on the state, and agents make inferences about aggregate past play conditional on their beliefs about the state.
@@
  • Discussion of continuum players. We follow the literature on global games by assuming that there is a continuum of atomless players. By the continuum law of large numbers, this implies that conditioned on the fundamental state $\theta$, all aggregate randomness in private signal realizations across the population washes out. This allows us to analyze the dynamics of aggregate play cleanly as the solution to a non-linear deterministic difference equation. In Appendix C we develop a finite-player version of our model with $N<+\infty$ players and show via standard concentration arguments of the empirical distribution of beliefs that limit play behaves similarly in the finite-player version of our model: for all $(\Sigma,\lambda_{0})$ and fundamental $\theta$, $\sup_{N\geq 2}\big{|}\lambda_{\infty}^{N}(\theta)-\lambda_{\infty}(\theta)\big{|}=0$ almost surely (see Theorem 2 in Appendix C).

### v2

@@ Abstract
We analyze inertial coordination games: dynamic coordination games with an endogenously changing state that depends on (i) a persistent fundamental players privately learn about over time; and (ii) past play. The speed of learning determines long-run equilibrium dynamics: the risk-dominant action is played in the limit if and only if learning is slow such that posterior precisions grow sub-quadratically. This generalizes results from static global games and endows them with a learning foundation. Conversely, when learning is fast such that posterior precisions grow super-quadratically, shocks can propagate and generate self-fulfilling spirals.
@@ 1   Introduction
Coordination games are at the heart of many dynamic economic phenomena. In such games, shocks to the fundamentals can sometimes propagate: as traders attack a currency peg and the central bank expands reserves to maintain the peg, more traders are incentivized to take advantage of this weakness which, in turn, further drains the central bank’s reserves, enticing even more attacks. But shocks can also fizzle out: traders might learn quickly that the central bank’s balance sheet is healthy so that the proportion of the population who are (incorrectly) pessimistic thins out rapidly. In this case, the downward spiral of an initial attack begetting further attacks eventually loses steam. [^1] Similar forces operate with bank runs (via bank liquidity), networked products (via product improvements), firm pricing (via sticky consumers), and deflationary spirals (via price and demand downturns). When do shocks propagate and when do they fizzle out?
To answer these questions, we study inertial coordination games: dynamic coordination games in which players repeatedly decide whether or not to take a risky action (e.g., attack a currency). The payoff of this risky action depends on both (i) a persistent fundamental component; and (ii) an endogeneous component which depends on past play (e.g., if more players have attacked a currency in the past, this depletes current reserves). Players privately learn over time and form inferences about the current state, which shapes their current play. This, in turn, propagates into the future by shaping the evolution of future states. Thus, our environment features two interconnected forces: private learning—in which dispersion in beliefs about fundamentals across the population shrinks with time—and intertemporal contagion, in which higher current play increases future states which incentivizes higher future play.
@@
Our main result (Theorem 1) develops a tight connection between the speed of learning and limit play: the risk-dominant action is played in the limit if and only if posterior precisions grow sub-quadratically. Sub-quadratic learning includes canonical learning environments such as (i) independent and identically distributed signals; and (ii) the receipt of a one-time signal. Super-quadratic learning includes social learning where in each period agents are randomly matched and sample their partner’s information sets. The tight connection we develop between learning speeds and limit play has simple but sharp economic implications. Slow (sub-quadratic) learning implies history independence: the initial shock to aggregate play as well the fine details of what agents learn in the interim is irrelevant for limit play. Conversely, when learning speeds are fast (super-quadratic), this induces history dependence such that initial shocks can be self-fulfilling: whether shocks propagate or not is jointly sensitive to both the fundamentals and the magnitude of the shock. [^2] See Section I of Krugman (1991) for a discussion of history-dependence and expectations in economics.
Our model of inertial coordination games imposes a lag between players’ actions and their effect on the time-varying state. Hence, when a player contemplates whether or not to play the risky action, his belief about the current state depends on his belief about past play. But that, in turn, depends on his belief about other players’ past beliefs which themselves are about play yet further back into the past. Thus, the path of past play (and beliefs)—even of the distant past—multiply forward to shape incentives in the present. [^3] Note that our environment is, in a loose sense, the opposite of coordination games with switching frictions (Frankel and Pauzner, 2000). In those models, switching frictions lead players to form beliefs about future play and hence future beliefs; by contrast, in our model, inertia leads players to form beliefs about past play and hence past beliefs. An important by-product of inertia is that it offers a simple yet novel means of obtaining unique predictions in coordination environments. At the same time, our model is motivated by—and consistent with—recent work documenting empirical facts about bank runs. Correia, Luck, and Verner (2024) find “most bank failures are the result of a deterioration in bank solvency” (in our model, the shock to initial play) rather than being purely self-fulfilling (predicted by static coordination games). Furthermore, they find that “depositors tend to be slow to react to an increased risk of bank failure” (in our model, the inertia). Hence, one of our contributions is to offer a simple dynamic model which reconciles empirical regularities about ‘real world’ coordination environments with the more abstract logic of learning and higher-order beliefs in coordination games. Further, our characterization of the connection between speed and limit play is invariant to the degree of inertia (duration of lag) and, as inertia vanishes, the equilibrium converges to an equilibrium of the model with purely contemporaneous incentives. [^4] This is made precise in Appendix B.
@@
In what sense does our model relate to the extant literature on coordination games? It turns out that static global games (Carlsson and Van Damme, 1993) are recovered exactly as a special case of our model in which players only receive a single signal at time $t=1$: since this learning process is sub-quadratic (posterior precisions are unchanged after $t=1$), the risk-dominant profile is—as in global games—played in the limit. This connection arises from observing that when there is no interim learning, best-responding to expected past play along each time step is equivalent in our model to the interim elimination of strictly dominated strategies in the static game. As we make precise in Online Appendix V, this uncovers a formal connection between “contagion” arguments over type spaces (which encode hierarchies of higher-order beliefs) and “contagion” over time (where the current state depends on past play). Thus, our results generalize the predictions of static global games to sub-quadratic learning and offer a simple dynamic learning foundation for the ‘vanishing noise limit’ commonly invoked in that literature.
@@
Finally, our model makes sharp qualitative predictions about the path of play (Proposition 3) whenever learning is slow enough such that the risk-dominant profile is played in the limit. In particular, transition dynamics toward limit play depend starkly on the precision of signals. When signals are precise, aggregate play exhibits a sudden phase transition from approximately all players playing the non-risk-dominant action (potentially for a long while) to approximately all players playing the risk-dominant action thereafter. Conversely, when signals are noisy, aggregate play exhibits a gradual transition as the measure of players playing risk-dominant action increases gradually. These differing transition regimes are driven by differences in time-varying heterogeneity in beliefs. The importance of heterogeneity has been recognized as a crucial determinant of equilibrium uniqueness in coordination games (Morris and Shin, 2006). Our results complement this insight by highlighting that the speed of learning—which drives the how the cross-sectional distribution of beliefs vary over time—is also a crucial determinant of the path of play. It also suggests that “spikes” in the time-series of aggregate actions (e.g., withdrawal decisions or adoption of networked goods) can be consistent with transition to limit equilibrium play and need not necessarily correspond to “equilibrium shifts” (Chwe, 2013; Morris and Yildiz, 2019).
@@
Related literature. Our paper directly relates to the large literature on global games. A broad takeaway from the literature on static global games (Carlsson and Van Damme, 1993; Kajii and Morris, 1997; Frankel, Morris, and Pauzner, 2003, among others) is that perturbing common knowledge by introducing a vanishing amount of private uncertainty often selects the risk dominant equilibrium. This approach has been widely deployed in a variety of economic applications to select among multiple equilibria (see, e.g., Morris and Shin (1998); Goldstein and Pauzner (2005) among many others). Our model nests the canonical global games model of Carlsson and Van Damme (1993) as a special case with a private signal at $t=1$ and no learning thereafter. The “contagion argument” in global games where players iteratively eliminate interim dominated strategies is sometimes viewed metaphorically or as a convenient way to select among equilibria. [^5] See, Morris (2002) and Chapter 2 of Spiegler (2024). Our simple dynamic environment with an endogeneously changing state thus offers an explicit dynamic foundation for the “contagion” logic of global games consistent with recently documented facts on bank runs (Correia, Luck, and Verner, 2024). Moreover, we substantially generalize the prediction of risk-dominant selection to the regime of sub-quadratic learning.
@@
Our work also relates closely to dynamic coordination games (Angeletos, Hellwig, and Pavan, 2007; Dasgupta, 2007; Dasgupta, Steiner, and Stewart, 2012, among others). In much of this literature, coordination is contemporaneous i.e., time-$t$ play matters for time-$t$ incentives. An important exception is the paper of Mathevet and Steiner (2013) who study finite-time horizon dynamic global games with non-contemporaneous coordination motives where the terminal outcome depends on both the state and path of play. [^7] See also the elegant paper of Dasgupta, Steiner, and Stewart (2012). Our models are non-nested and deliver complementary insights. In our model, the state is time-varying and endogenous to past play so players’ payoffs depend on interim play which in turn drives the state-evolution. In Mathevet and Steiner (2013), ex-post payoffs depend on a binary outcome which is determined by a persistent fundamental and the path of aggregate play. This mapping from paths to a binary outcome allows Mathevet and Steiner (2013) to tractably handle a wide class of payoffs. By contrast, our payoffs are more specific, but evolve endogeneously and take on a continuum of values. We exploit the structure of our environment to derive sharper implications of learning speeds on limit outcomes. [^8] The main result of Mathevet and Steiner (2013) is that aggregate play at the critical state at which outcomes hinge only depends on payoffs; by contrast, we show that limit aggregate play at any state depends sharply on the speed of learning. Note also that Mathevet and Steiner (2013)’s definition of “fast learning” takes the precision of each signal arbitrarily large, which is sufficient for equilibrium existence. This is different from our notion which requires the (weaker) condition of super-quadratic growth in posterior precisions, and where limit equilibrium play always converges.
Our model can be viewed as the conceptual opposite of models which impose switching frictions to obtain unique equilibrium predictions (Frankel and Pauzner, 2000). [^9] See also Frankel, Morris, and Pauzner (2003) and more recent work in dynamic information design by Koh, Sanguanmoo, and Uzui (2024). Angeletos and Lian (2016) survey work in macroeconomics applying switching frictions to obtain equilibrium selection. A key feature of our model is that players can frictionlessly change their actions, but there is a (potentially small) lag between players’ actions and transmission to the state. This lag links incentives intertemporally: when a player contemplates whether or not to play the risky action, he must form beliefs about past play and so beliefs about even the distant past can multiply forward to shape present incentives. Thus, our model can be loosely viewed as the conceptual counterpart of models with switching frictions which require that players form beliefs about future play (Frankel and Pauzner, 2000). Unlike such models which predict that the risk-dominant action will be selected, our model offers a more nuanced unique prediction which hinges on the speed of private learning.
@@
Finally, our results are connected to the literature on learning in coordination games. An interpretation of our model is of best-response dynamics with Bayesian learning about a fundamental state or past play. To our knowledge, this is novel to the extant literature on learning in games which assumes that past play is perfectly observed and there is no fundamental uncertainty. [^10] See Fudenberg and Levine (1998) for a survey. Our paper is related to an important paper of Crawford (1995) who studies adaptive dynamics in complete information coordination games by modelling player strategies as a linear adjustment rule to be estimated; this is similar to the interpretation of players in our model as best-responding to past play. Thus, a contribution of our paper is to analyze how Bayesian learning about fundamental uncertainty interacts with adaptive learning about strategic uncertainty (Camerer and Hua Ho, 1999). [^11] Camerer and Hua Ho (1999) develop a model of learning in complete information games which features both modes of learning; by contrast, in the best-response interpretation of our model, different modes of learning correspond to different kinds of uncertainty: ‘fundamental uncertainty’ is resolved via Bayesian learning but ‘strategic uncertainty’ is resolved via adaptive or reinforcement learning.
@@ 2   Model
There is an unknown and persistent fundamental $\theta\in\mathbb{R}$. For each $t\in\mathcal{T}$, let $\theta_{t}$ be the state at time $t$ and for each $t>1$, [^13] For any $\alpha\in(0,1)$, analogs of our results apply for $\theta_{t}=\alpha\theta+(1-\alpha)\lambda_{t-1}$; see Online Appendix VI.
@@
Our law of motion for the endogenous state is simple. Nonetheless, we think it well-approximates a range of environments such as currency attacks or bank runs in which the failure probability depends on a single sufficient statistic: present reserves. This, in turn, is simply the sum of its initial reserves less the ‘current’ short/withdrawal positions. For a currency attack where the risky action ($a=1$) is to attack, initial reserves are $-\theta$ and the current short position is $\lambda_{t-1}$; for a bank run the risky action ($a=1$) is to deposit, initial reserves are $\theta$, and the current withdrawal position is $1-\lambda_{t-1}$. [^14] As is standard in global games, we have assumed that $\theta$ is supported on the real line, but it can be appropriately shifted in space and truncated so that the sign fits the economic phenomena in question.
@@
$\lambda_{0}$ is common knowledge but interim payoffs as well as $(\lambda_{t})_{t\geq 1}$ are unobserved and inferred via beliefs about the state $\theta$ [^15] An identical assumption is made in the context of dynamic global games by Angeletos, Hellwig, and Pavan (2007). Our model is equivalent to one in which agents observe noisy signals of past play $(\lambda_{t})_{t\geq 1}$, or both signals about the state as well as past play—we formalize this an equivalence in Appendix C and discuss this below.
Interpreting initial play. We interpret $\lambda_{0}\in[0,1]$ as a shock. This is equivalent to a one-off and transitory shock to the cost at time $1$. [^16] More explicitly, suppose there is zero initial play but there is an initial shock to costs of playing the risky action. Costs are potentially time-varying $c_{1},c_{2},\ldots$ with $c:=c_{2}=c_{3}=c_{4}=\ldots$ and $c_{1}\neq c$. Then, the dynamics of this model with zero initial play is equivalent to the model with $\lambda_{0}=c-c_{1}$. The impact of this shock is equivalent to that of non-zero initial play: positive shocks which incentive the risky action corresponds to high $\lambda_{0}$ and vice versa. In the language of macroeconomics, this corresponds to an unexpected shock (“MIT shock”) and we analyze the dynamics of aggregate play following the shock. In the language of fictitious play, this corresponds to a shock to initial play and we analyze how the speed of learning shapes the limit action profile.
@@
Limits and Learning. Let the learning process $\Sigma:=(\sigma_{t}^{2})_{t\in\mathcal{T}}$ be a strictly positive sequence of signal variances determining the speed of learning. Note that instead of $\Sigma$, we may equivalently directly specify a nonnegative, increasing sequence of posterior precisions $(\eta_{t}^{-2})_{t}$, since such a sequence is induced by $\sigma_{t}^{-2}:=\eta_{t}^{-2}-\eta_{t-1}^{-2}$. When convenient, we will therefore take $(\eta_{t}^{-2})_{t}$ as a primitive of our model.
@@
We discuss several aspects of our model:
  • State transition & non-contemporaneous incentives: Our model is one in which the state at time-$t$ ($\theta_{t}$) is shaped by play at time-$t-1$ ($\lambda_{t-1}$). A few remarks are in order. First, we think this captures an important element of dynamic coordination phenomena. For instance, in currency crises the susceptibility of the current regime depends on the short position. Note also that our model exhibits path dependence—though $\theta_{t}$ depends directly only on $\lambda_{t-1}$, it also depends indirectly on the full path of past play $(\lambda_{s})_{s<t-1}$. [^20] This is broadly analogous to how an AR$(1)$ process depends indirectly on realizations of the process which are far into the past (although our setting is non-stationary). An alternate interpretation is that our model is a form of best-response dynamics. Of course, some care is required with this interpretation since past play is unobserved: players are thus best-responding to their best guess of past play. Experimental evidence in (complete information) coordination games find that the vast majority (96%) of play constitute myopic best-responses (Mäs and Nax, 2016). This allows our model to make rich and nuanced predictions on how the speed of learning shapes equilibrium selection. By contrast, a model with exact contemporaneous incentives cannot capture this—the risk-dominant equilibria is always selected. [^21] In Appendix B we show that the unique belief threshold is $c-1/2$ so the risk-dominant action is always selected in the limit. Second, our main result on learning regimes (Theorem 1) holds for vanishing inertia: if we fixed the speed of learning as a function of calendar time but shrink the spacing between time periods, the sub/super-quadratic threshold for risk/non-risk dominance continues to hold. This is in the same spirit as the ‘frictionless limit’ of Frankel and Pauzner (2000) where switching frictions vanish. Third, as the degree of inertia vanishes, the (unique) equilibrium of our inertial coordination game converges to an equilibrium of the coordination game with contemporaneous incentives. [^22] That is, with utility function $u(1,\theta,\lambda_{t})=\theta+\lambda_{t}$. Proposition 4 in Appendix B formalizes this. This means that the speed of learning relative to the degree of inertia is what is important for obtaining richer predictions about equilibrium outcomes.
@@
  • Observability. We have assumed agents do not observe aggregate past play. [^23] Aggregate past play is an injective function of the state, so if agents were to observe this, they would immediately learn the state. However, our model is equivalent to one in which agents receive private signals both about the state, and about aggregate past play. [^24] For instance, see Dasgupta (2007); Trevino (2020). Explicitly, suppose that agent $i\in[0,1]$, in addition to observing $x_{it}\sim N(\theta,\sigma^{2}_{t})$, additionally observes an independent signal distributed $y_{it}\sim N(\Phi^{-1}(\lambda_{t-1}),\tau_{t}^{2})$; this signal structure was introduced by Dasgupta (2007) and used recently by Trevino (2020) to study financial contagion. Then, we can show by induction that equilibrium behavior is equivalent to a model in which players receive a more precise signal about $\theta$ and no information about aggregate past play; Appendix C formalizes this connection. In light of this observation, we load all learning on the state, and agents make inferences about aggregate past play conditional on their beliefs about the state.
@@
  • Continuum of players. We follow the literature on global games by assuming that there is a continuum of atomless players. By the continuum law of large numbers, this implies that conditioned on the fundamental state $\theta$, all aggregate randomness in private signal realizations across the population washes out. This allows us to analyze the dynamics of aggregate play cleanly as the solution to a non-linear deterministic difference equation. In Online Appendix IV we develop a finite-player version of our model with $N<+\infty$ players and show via standard concentration arguments of the empirical distribution of beliefs that limit play behaves similarly in the finite-player version of our model (see Theorem 2 in Online Appendix IV).